OpenStudy (anonymous):

We have 2.8 meters of rope. How many centimeters is that? 2.8 28.0 280 2800

OpenStudy (anonymous):

2.8 x 100= 280

OpenStudy (ash2326):

@Lissaa_bih 1 meter=100 centimeters so 2.8 meters= 2.8*100 centimeters Can you tell which is the correct option?

OpenStudy (ash2326):

@Lissaa_bih where do you have doubt?

OpenStudy (anonymous):

the part that u had put 100 centimeters , so 2.8 meters = 2.8*100 centi .

OpenStudy (ash2326):

Ok, 1 meter = 100 centimeters multiply both side by 2.8 so 2.8*1 meter=2.8*100 centimeters or 2.8 meters=2.8*100 centimeters

OpenStudy (ash2326):

@Lissaa_bih do you get it?

OpenStudy (anonymous):

so i multiply 100*2.8 twiice?

OpenStudy (ash2326):

no just once 100 *2.8

OpenStudy (anonymous):

@ash2326 280 is my answer ?!?!?!?

OpenStudy (ash2326):

correct:)

OpenStudy (anonymous):

its not that complicated dont make it more than it is.. 100 cent = 1 meter. dont break ur head

OpenStudy (ash2326):

@Vikt0r Please don't behave rudely. The user is trying to learn, everyone has a different pace of learning. Thanks

OpenStudy (anonymous):

im dislexic @Vikt0r sorry im not good at math , k thanks & @ash2326 thanks for helping and understanding .

OpenStudy (anonymous):

lol.... i have severe autism.

OpenStudy (ash2326):

@Lissaa_bih you don't have to be sorry. It's completely ok. I'm sure you can be good at Math:)

OpenStudy (anonymous):

@ash2326 thank you(:

OpenStudy (ash2326):

Glad to help :)

OpenStudy (anonymous):

& @Vikt0r im sorry to hear /;